# Expressivity of Neural Networks with Random Weights and Learned Biases

### v1

@@ Appendix D Supplementary Lemmas
It is easy to see that $\hat{z}_{t}=Ar_{t}$ for all $0\leq t\leq T$. It follows that $\hat{y}_{t}=CAr_{t}\>\forall t$. Thus, this RNN approximates the original partially observed dynamical system.
[…]
It remains to treat the initial values of the RNN, $r_{0}$. These are defined implicitly as $Ar_{0}=z_{0},z_{0}\in U_{z}$. We will show that we can span all $z_{0}\in U_{z}$ by selecting $r_{0}$ from a compact set. By assumption $A$ is full rank so $\mathrm{col}(A)=\mathbb{R}^{s}$ and, thus, we can find a subset of the columns of $A$ that form a basis for $\mathbb{R}^{s}$. Let these columns be $\{A_{:i}\}_{i\in\nu}$. Construct a matrix $A_{\nu}$ by deleting the columns whose indices are not contained in $\nu$. Because the columns of $A_{\nu}$ are a basis the matrix is invertible. Define $U_{r}^{(0)}=\{r:r_{i}=\tilde{r}_{i}\>\mathrm{if}\>i\in\nu,r_{i}=0\>\mathrm{if}\>i\not\in\nu,\tilde{r}=A_{\nu}^{-1}z_{0},z_{0}\in U_{z}\}$. For each $z_{0}$ in $U_{z}$ we have an initial condition $r_{0}\in U_{r}^{(0)}$ and, moreover, $r_{0}$ is a continuous mapping of $z_{0}$ (via $A_{\nu}^{-1}$). Since $z_{0}$ is taken from a compact set $U_{r}^{(0)}$ must also be compact.
Finally, define $U_{r}=U_{r}^{(0)}\cup U_{r}^{(1)}$. From the above this contains all hidden states of the RNN for the finite time trajectories of interest from any initial condition. As a union of two compact sets it is also compact. We will use $U_{r}$ in Lemma 2. We note that the size of the set will likely depend, via the dependency on $A$, on the function being approximated and on the neural network hidden layer size $n$.
Lemma 4.
Eq.B.7 holds under the assumptions of Lemma 1.
Proof.
This follows by adding and subtracting $A_{ij}^{\mathcal{M}}\phi(B_{j:}x+b_{j})$ inside the summations in the norm and the expanded neural network expression, and then successively bounding things using the assumptions on the domain and the activation function:
$||y_{m}^{\mathcal{M}}-y_{n}||$ $\leq\sum_{i=1}^{l}\sum_{j=1}^{n}\Big{|}A_{ij}^{\mathcal{M}}\big{[}\phi(B_{j:}^{\mathcal{M}}x+b_{j}^{\mathcal{M}})-\phi(B_{j:}x+b_{j})\big{]}-(A_{ij}-A_{ij}^{\mathcal{M}})\phi(B_{j:}x+b_{j})\Big{|}$ (D.13)
$\leq\sum_{i=1}^{l}\sum_{j=1}^{n}\Big{[}\bar{\gamma}K_{\phi}\big{|}(B_{j:}^{\mathcal{M}}-B_{j:})x+b_{j}^{\mathcal{M}}-b_{j}\big{|}+\varepsilon M_{\phi}\Big{]}$ (D.14)
$\leq nl\big{[}\bar{\gamma}K_{\phi}(\varepsilon M_{x}+\varepsilon)+\varepsilon M_{\phi}\big{]}=nl\big{[}\bar{\gamma}K_{\phi}(M_{x}+1)+M_{\phi}\big{]}\varepsilon.$ (D.15)
Lemma 5.
Eq.C.6 holds under the assumptions of Lemma 2.
Proof.
We have:
$||r_{0}^{\mathcal{M}}-r_{0}||$ $=0$ (D.16)
$||r_{1}^{\mathcal{M}}-r_{1}||$ $\leq\alpha||r_{0}^{\mathcal{M}}-r_{0}||+\beta K_{\phi}\big{[}\bar{\gamma}n||r_{0}^{\mathcal{M}}-r_{0}||+\varepsilon n(R_{r}+R_{x}+1)\big{]}$ (D.17)
$=\beta K_{\phi}\bar{\gamma}n\tilde{R}\varepsilon\leq 1,$ (D.18)
where we used that $r_{0}=r_{0}^{\mathcal{M}}$ and thus both are contained in $U_{r}$ so that the arguments of $\phi$ are contained in $\tilde{U}$, allowing us to use the Lipschitz result and constant $K_{\phi}$. We note that the final inequality follows from the choice of $\varepsilon$ and means that, since $r_{1}\in U_{r}$, $r_{1}^{\mathcal{M}}=r_{1}+r^{\prime}$ with $||r^{\prime}||\leq 1$. This means that for the next step forward in time the arguments of $\phi$ will once again be in $\tilde{U}$ so that the Lipschitz conditions will continue to hold. Using $t=1$ as a base case we make the induction hypothesis:
$||r_{t}^{\mathcal{M}}-r_{t}||$ $\leq n\beta K_{\phi}\tilde{R}\sum_{t=0}^{t-1}(\alpha+\beta K_{\phi}n\bar{\gamma})^{t}\varepsilon$ (D.19)
$||r_{t}^{\mathcal{M}}-r_{t}||$ $\leq 1.$ (D.20)
Using the induction hypothesis it is straightforward to prove the induction step, and to show that $||r_{t+1}^{\mathcal{M}}-r_{t+1}||$ remains smaller than $1$ and thus that the arguments of $\phi$ remain in $\tilde{U}$. Proving the induction step and then bounding the output matrix elements by $\bar{\gamma}$ yields the result.

### v3

@@ B.4 Supplementary Lemmas
This function is clearly continuous and satisfies $Ar_{0}=\sum_{i\in\nu}A_{:i}{r_{0}}_{i}=\sum_{i\in\nu}A_{:i}{A_{\nu}^{-1}}_{i:}z_{0}=A_{\nu}A_{\nu}^{-1}z_{0}=z_{0}$. If we then define $r_{t}=-\alpha r_{t-1}+\phi(WAr_{t-1}+Bx_{t-1}+b)$ for all $1<t\leq T$ we see that $r_{t}$ is simply the state variable for an RNN of the style of Eq.2, and that it satisfies $\hat{z}_{t}=Ar_{t}$ for all $0\leq t\leq T$. It follows that $\hat{y}_{t}=CAr_{t}\>\forall t$. Thus, this RNN approximates the original partially observed dynamical system in the sense described in section §2.2.
Appendix C Methods for Numerical Results Section
C.1 Methods for Section 3.1
In this section all networks were single hidden layer FNNs. For figure 1A, networks with widths of $64$, $128$, $256$, $512$, $1024$, $2048$, $4096$, $8192$, $16384$, and $32768$ were trained, with the width being indicated on the $x$-axis. For figures 1B and 1C, a width of $3.2\times 10^{4}$ was used.
For figure 1A, all networks were trained on FashionMNIST, with $5\times 10^{4}$ training samples and $10^{4}$ test samples, using ADAM with a learning rate of $0.01$. Training was run for $20$ epochs with a batch size of $512$.
Xavier uniform initialization with a gain of $1.0$ was used for the fully trained networks, while the frozen weights for the bias-only networks were sampled from either a uniform distribution on $[-0.1,0.1]$ or from a zero-mean Gaussian with standard deviation $\frac{1}{\sqrt{d}}$, where $d$ is the input dimension.
For Fig.E.3 a network with $3.2\times 10^{4}$ units was trained on $2\times 10^{5}$ random samples from a concatenation of the $8$ MNIST style datasets used in Figure 1. Weights were initialized to Gaussians with mean 00 and SD $\frac{1}{\sqrt{d}}$, biases were initialized uniformly on $[-0.01,0.01]$, and ADAM with a learning rate of $1\times 10^{-5}$, and batch size $512$ was used to optimized a cross entropy loss for $15$ epochs. The network reached test accuracies of 0.9924 (Ethiopic) 0.8624 (Fashion), 0.8883 (Kannada), 0.0000 (KMNIST), 0.9354 (MNIST), 0.9995 (Nko), 0.9979 (Osmanya), and 0.9910 (Vai). Thus, it failed to learn KMNIST, which we hypothesize to be due to interference with the other datasets. In this figure and in Fig.1 we used the Kmeans from Scikit-Learn with 20 clusters. Before clustering/plotting we normalized the variances by dividing each $8$ dimensional vector of unit variances by the maximal value plus a small constant to avoid division by zero.
C.2 Methods for Section 3.2
In this section all networks were single hidden layer FNNs with $10^{4}$ ReLU units. Weights were each initialized to a Gaussian with mean 00 and standard deviation of $1/28$–the inverse square root of the input dimension–and were left unchanged during training.
Both bias and mask networks were trained on MNIST, with $5\times 10^{4}$ training samples and $10^{4}$ test samples, using ADAM with a learning rate of $0.01$. Trained parameters were each initialized uniformly on $[-0.01,0.01]$ (for the mask learned networks the bias vector was initialized in this way and left untrained), and training was run for $30$ epochs with a batch size of $512$. For bias learning, the trained parameters were the $10^{4}$ element bias vector, $b$; for mask learning, trained parameters were a $10^{4}$ element vector of gains $g$.
The output of the $i^{th}$ unit in the mask-trained network was $\mathrm{ReLU}\big{(}W_{i:}\cdot x+b\big{)}\mathrm{Sigmoid}\big{(}\frac{g_{i}}{\tau}\big{)}$. Here, $x$ was the input, flattened, MNIST image, $W$ the random hidden weight matrix and $\tau$ an annealing parameter. Starting from $\tau_{0}=1$, at each epoch $\tau$ was decreased so that at epoch $k$ $\tau_{k}=c\tau_{k-1}$, with $c$ chosen such that at the final epoch $\tau_{30}=0.001$. This scheme was chosen so that by the end of training the sigmoid functions would be almost step functions, well-approximating the desired binary mask. We found that this worked better than immediately setting $\tau=0.001$. For testing, units were evaluated with the $\mathrm{Sigmoid}$ factor binarized ($\tau=\infty$).
The null model used in Fig.2.B was calculated by drawing $10^{4}$ samples from two, independent, Bernoulli random variables whose probability of being ON (drawing a $1$) were equal to the mean fraction of ON units in the bias and mask-trained networks respectively.
C.3 Methods for Section 3.3
C.3.1 Cosine generation (Fig. 3A-B)
We used $N=200$ hidden recurrent units with ReLU activations. Biases were initialized from $\mathcal{U}(0,1)$. The recurrent weights were sampled from $\mathcal{N}(0,N^{-1})$ and the output weights from $\mathcal{N}(0,N^{-2})$. The target period of the cosine was $25$ and the total duration to generate was $125$. Learning rates were $0.1$ for bias learning and $0.001$ for the fully-trained network; other parameters for the Adam optimizer were left at their default values in Pytorch.
C.3.2 Van der Pol oscillator (Fig. 3C-D)
We used $675$ hidden ReLUs for the bias-learning network and $25$ for the fully-trained network. The recurrent weights were sampled from $\mathcal{N}(0,gN^{-1})$, where $g$ is the “gain”; other weights were initialized as in the previous subsection. The Van der Pol oscillator obeyed $\ddot{x}=\mu(1-x^{2})\dot{x}-x$, for $\mu=2$. Learning rates were $0.1$ for bias learning and $10^{-4}$ for the fully-trained network and other optimization parameters were as above.
C.4 Methods for Section 3.4
C.4.1 RNN architecture and hyperparameters
The architecture is a single-layer one-dimensional input/output recurrent neural network with ReLU activations. The training optimizer was Adam with a learning rate of $0.001$ and a weight decay value of $0.1$.
C.4.2 Lorenz attractor
The equations used to generate the partially-observed Lorenz attractor were the following:
$\frac{dx}{dt}=\sigma(y-x),\quad\frac{dy}{dt}=x(\rho-z)-y,\quad\frac{dz}{dt}=xy-\beta z$
where the initial point is $(0,1,0)$ and $\sigma$, $\rho$, and $\beta$ are $10$, $28$, and $\frac{8}{3}$, respectively. The trajectory was generated using Euler’s method with a step size of $0.01$.
C.5 Methods for Section 3.5
We used $N=1,\!024$ hidden ReLU units in a vanilla RNN (Eq. 2 with $\alpha=0$ and $\beta=1$). The input to the network was the 2D Cartesian position of the targets, which were 0.07 m away from the center of the workspace. The point-mass arm obeyed a discrete-time version of the following dynamics
$\frac{dx}{dt}=\dot{x}(t),\quad\frac{d\dot{x}}{dt}=\frac{f(t)}{m},\quad\frac{df}{dt}=\frac{u(t)-f(t)}{\tau_{f}}$
where $x$ is the 2D arm position, $\dot{x}$ its velocity and $f$ the applied force. The initial conditions were $x(0)=\dot{x}(0)=f(0)=0$, i.e., the arm was initially at rest. The force was obtained from an exponential filtering of the controls $u(t)$ with time constant $\tau_{f}=0.04$ s. The network generated the controls $u(t)$ via a linear readout of its activity: $u=Cr$. The input ($B$), recurrent ($W$) and output ($C$) matrices were initialized as: $B_{ij}\sim\mathcal{U}(-1/\sqrt{2},1/\sqrt{2})$, $W_{ij}\sim\mathcal{N}(0,1.5625/N)$, and $C_{ij}\sim\mathcal{N}(0,0.5/N)$. The loss function was
$L=\sum_{k=0}^{K-1}\frac{\|x^{(k)}_{T}-d_{k}\|^{2}}{\delta_{p}^{2}}+\gamma_{v}\frac{\|\dot{x}^{(k)}_{T}\|^{2}}{\delta_{v}^{2}}+\gamma_{f}\frac{\|f^{(k)}_{T}\|^{2}}{\delta_{f}^{2}},$ (31)
where $d_{k}=D[\cos(2\pi k/K),\sin(2\pi k/K)]^{\top}$, $k=0,\ldots,K-1$, represent the position of the $K=6$ peripheral targets at a distance $D=0.07$ m from the center target. Here, $x^{(k)}_{T}$, $\dot{x}^{(k)}_{T}$ and $f^{(k)}_{T}$ are the final position, velocity and force for the $k$th target ($T=1\mathrm{s}/0.01\mathrm{s}=100$, where $0.01$ s was our integration time step). Therefore, the objective was to reach the target at the end of the trial (first term) with near-zero velocity (second term) and force (third term). Parameters $\delta_{p}=0.01$, $\delta_{v}=0.02$ and $\delta_{f}=0.08$ were used to rescale the position, velocity and acceleration terms. Hyperparameters $\gamma_{v}=0.2$, and $\gamma_{f}=0.04$ controlled the relative weight of the velocity and force costs with respect to the position loss. The learning rate of the standard Adam optimizer was set to $3\times 10^{-3}$ and training was stopped when a loss of $5\times 10^{-3}$ was reached.
Appendix D Training Time
We ran a brief experiment on training time, which we report as $mean\pm 1SD$ over 5 seeds. Training a fully-trained single-layer MLP with $10^{4}$ hidden units on MNIST for 5 epochs takes $69.17\pm 0.22$ seconds. Training only biases for the same task and architecture takes $61.51\pm 0.92$ seconds. Thus, for matched layer widths training biases with standard Pytorch implementations provides a slight advantage. Note that one can maintain performance and get faster training by simply reducing the width of the fully-trained network: a 100 hidden unit fully trained net takes $37.77\pm 0.04$ on the same test.
Appendix E Supplementary Figures
Figure E.1: Proof Intuition. Step 1. Consider a smooth, partially observed Dynamical System (DS) on an invariant set, $\Omega$ (or a function $y(x)$ on a compact set $\Omega$). Step 2. Let $0<t\leq T$. Approximate DS on $\Omega$ using RNN $\mathcal{R}_{1}$ (or approximate $y$ on $\Omega$ using FNN $\mathcal{N}_{1}$). Step 3. Randomly initialize a larger RNN, $\mathcal{R}_{2}$ (or FNN $\mathcal{N}_{2}$). Check for a sub-network of units with parameters matching $\mathcal{R}_{1}$ ($\mathcal{N}_{1}$). Step 4. Adjust biases outside the sub-network to be very negative–thus shutting off the units outside the sub-network–and those inside the sub-network to match those of $\mathcal{R}_{1}$ ($\mathcal{N}_{1}$).
Figure E.2: Supplementary Experiments. A. As in Fig.1.A except with $x$ and $y$-axes log-scaled. B. As in Fig.1.A except $x$-axis is layer width. C. Performance on MNIST (y-axis) of bias learned MLPs as function of distribution from which each individual weight is initialized (x-axis). Orange: uniform on $[-\gamma,\gamma]$; Blue: 00 mean Gaussian with standard deviation $\sigma$; Green: mixture of Gaussians centred at ±$\pm\mu$ each with standard deviation $0.015$. Layer width is $10,000$ units. D. As in Fig.4.A except $x$-axis is layer width. E. Performance ($y$-axis) of fully trained versus bias learned (uniform initialized weights) MLPs on fashion MNIST as a function of depth ($x$-axis). Layer width is 2048 units for both learning types. All plots: plotted are means and $\pm$±SD over $5$ random seeds.
Figure E.3: Supplementary for Figure 1. Figure 1.C but for a single fully-trained network fit to a meta-dataset containing the $8$ MNIST style datasets used in Fig.1. The bottom row is cluster index. We observe task selectivity that is qualitatively similar to bias-learning. See §C.1 for methods.
Figure E.4: Supplementary for Figure 2. A. Decrease in bias learning test accuracy as the task variance threshold, above which units are ON, is increased. $x$-axis is number of ON units; $y$-axis is test accuracy on MNIST. Threshold values are those used in Fig.2.B. Plotted is mean and SD over $3$ training runs. B. Probability that a given unit is ON/OFF for versions of a network with the same weights trained by bias versus mask learning. Left group of bars is probability of the unit being ON for bias learning given that it is ON for mask learning, middle is the same but for OFF instead of ON, right is the probability of either match occurring (OFF or ON in both training paradigms). Grey is a null model (see section C.2 for details); orange is the bias-mask comparison; black is the same comparison but between two mask-learning training runs. C. Scatter plot of hidden unit variances taking only units that are ON for both mask and bias learning; bias-trained on x-axis and mask-trained on y-axis. Error bars are over $5$ training runs. We observe a mean correlation coefficient of $0.5030\pm 0.0087$. We note that the correlation between two mask-learning runs on the same random weights is $0.9999\pm 0.0000$. Correlation coefficients are over $5$ samples; plot data is from one, representative, pair of bias/mask networks.
Figure E.5: Supplementary for Figure 5. Same figure as Fig. 5 in the main text, but for fully-trained networks. The architecture and initial parameters (seeds) were the same as for the bias-learning networks (see Methods). The only change was that the learning rate was $10^{3}$ times smaller than that for bias learning. For these matched parameter initializations (in particular, given the high variance for the recurrent weights), the solutions found by the fully-trained network had more temporal variations compared to bias learning.